OpenStudy (anonymous):

Imagine you are teaching a fellow student how to solve: 4 + 2(x - 1) = 5x - (x + 10) In your own words, explain the process for solving this equation. Please include your work at each step of the process along with the final answer

OpenStudy (anonymous):

Easy squeezy Lemon Peezy. Haha. Um okay so...

OpenStudy (anonymous):

4 + 2(x - 1) = 5x - (x + 10) 4 + 2x - 2 = 5x - x - 10 2x - 5x + x = - 4 + 2 - 10 -2x = -12 x = -12/-2 x = 6

OpenStudy (anonymous):

I for got all the proper terms for it but multiply two times x and two times negative 1 then the sum of that answer plus 4. That negative on the second half has an invisible one so negative one times x plus 10 and that x becomes a negative 1x so negative 1x plus 5x is 6 x so x equals six or if she wants you to write it out with numbers the one above me is right too.

OpenStudy (anonymous):

*forgot

OpenStudy (anonymous):

Now idk if I confused myself or not.. Bleh go with the top!